The vapour pressure of water at 323k is 92mm Hg.18.1g of urea dissolved in 100g of water.calculate the lowering of vapour pressure?


1
Expert's answer
2021-01-04T03:47:07-0500

moles urea = 18.1 / 60.06 = 0.3014 mol

moles water = 100 / 18.015 = 5.551 mol

total moles = 0.3014 + 5.551 = 5.8524 mol

mole fraction of solvent = 5.551 / 5.8524 = 0.9485 (94.85%)

Raoult's law:

P solution = X solvent x P0solvent

P solution = 0.9485 x 92 mm Hg = 87.3 mm Hg.